Metlock financial income for Lake Inc. is $320,000, and its taxable income is $90,000 for 2018. Its only temporary difference at the end of the period relates to a $70,000 difference due to excess depreciation for tax purposes. If the tax rate is 38% for all periods, compute the amount of income tax expense to report in 2018. No deferred income taxes existed at the beginning of the year.

Income tax expense $
Answer #1
Income tax expense is $60,800
Dear Student
Thank you for using homeworklib
Please find below the answer
Statement showing Computations
Particulars Amount
Taxable income              90,000.00
Tax rate 38%
Current tax expense =90000*38%              34,200.00
Futuretaxable amounts              70,000.00
Tax rate 38%
Deferred tax expense = 70,000*38%              26,600.00
Income tax expense = 34200 + 26600              60,800.00